A 2-m long wire with a mass of 70 g is under tension. A transverse wave for which the frequency is 570 Hz, the wavelength is 0.6 m, and the amplitude is 4.6 mm is propagating on the wire. The maximum transverse accekeration of a point on a wire is closest to?
Answer #1

from the given data

angular frequency, w = 2*pi*f

= 2*pi*570 rad/s


maximum transverse acceleration, a_max = A*w^2

= (4.6*10^-3*(2*pi*570)^2

= 5.9*10^4 m/s^2